OpenStudy (anonymous):

what is the equation of a line that has a slope of 1/2 and passes through the point (2-3)?

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

im lazy do it for me??

OpenStudy (anonymous):

\[\huge {\color{Blue} WELCOME.TO.OPESTUDY!!!!!!!!!!!!!}\]

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

y1= y in the point x1= x in the point, just substitute it.

OpenStudy (anonymous):

it will be in a form of y=mx+b